# A Comprehensive Overview of Large Language Models

### 27 Dec 2023 v7

@@ Abstract
Figure 1: The trends in the number of LLM models introduced over the years.
@@ I Introduction
Additional to better generalization and domain adaptation, LLMs appear to have emergent abilities, such as reasoning, planning, decision-making, in-context learning, answering in zero-shot settings, etc. These abilities are known to be acquired by them due to their gigantic scale even when the pre-trained LLMs are not trained specifically to possess these attributes [22, 23, 24]. Such abilities have led LLMs widely adopted in diverse settings including, multi-modal, robotics, tool manipulation, question answering, autonomous agents, etc. Various improvements have also been suggested in these areas either by task-specific training [25, 26, 27, 28, 29, 30, 31] or better prompting [32].
Figure 3: A broader overview of LLMs, dividing LLMs into five branches: 1. Training 2. Inference 3. Evaluation 4. Applications 5. Challenges
The LLMs abilities to solve diverse tasks with human-level performance come at a cost of slow training and inference, extensive hardware requirements, and higher running costs. Such requirements have limited their adoption and opened up opportunities to devise better architectures [15, 33, 34, 35] and training strategies [36, 37, 21, 38, 39, 40, 41]. Parameter efficient tuning [38, 41, 40], pruning, quantization, knowledge distillation, and context length interpolation [42, 43, 44, 45] among others are some of the methods widely studied for efficient LLM utilization.
@@
  • We present a survey on the developments in LLM research with the specific aim of providing a concise yet comprehensive overview of the direction.
@@
  • Besides paying special attention to the chronological order of LLMs throughout the article, we also summarize major findings of the popular contributions and provide detailed discussion on the key design and development aspects of LLMs to help practitioners to effectively leverage this technology.
  • In this self-contained article, we cover a range of concepts to comprehend the general direction of LLMs comprehensively, including background, pre-training, fine-tuning, robotics, multi-modal LLMs, augmented LLMs, datasets, evaluation, etc.
We loosely follow the existing terminologies to ensure providing a more standardized outlook of this research direction. For instance, following [46], our survey discusses pre-trained LLMs with 10B parameters or more. We refer the readers interested in smaller pre-trained models to [47, 48, 49].
The organization of this paper is as follows. Section II discusses the background of LLMs. Section III focuses on LLMs overview, architectures, training pipelines and strategies, and utilization in different aspects. Section IV presents the key findings derived from each LLM. Section V highlights the configuration and parameters that play a crucial role in the functioning of these models. Summary and discussions are presented in section VIII. The LLM training and evaluation, datasets and benchmarks are discussed in section VI, followed by challenges and future directions and conclusion in sections IX and X, respectively.
@@ II Background
We provide the relevant background to understand the fundamentals related to LLMs in this section. Aligned with our objective of providing a comprehensive overview of this direction, this section offers a comprehensive yet concise outline of the basic concepts. We focus more on the intuitive aspects and refer the readers interested in details to the original works.
@@ II-A Tokenization
LLMs are trained on text to predict text, and similar to other natural language processing systems, they use tokenization [55] as the essential preprocessing step. It aims to parse the text into non-decomposing units called tokens. Tokens can be characters, subwords [56], symbols [57], or words, depending on the size and type of the model. Some of the commonly used tokenization schemes in LLMs are briefed here. Readers are encouraged to refer to [58] for a detailed survey.
II-A1 WordPiece [59]
It was introduced in [59] as a novel text segmentation technique for Japanese and Korean languages to improve the language model for voice search systems. WordPiece selects tokens that increase the likelihood of an n-gram-based language model trained on the vocabulary composed of tokens.
II-A2 BPE [57]
Byte Pair Encoding (BPE) has its origin in compression algorithms. It is an iterative process of generating tokens where pairs of adjacent symbols are replaced by a new symbol, and the occurrences of the most occurring symbols in the input text are merged.
II-A3 UnigramLM [56]
In this tokenization, a simple unigram language model (LM) is trained using an initial vocabulary of subword units. The vocabulary is pruned iteratively by removing the lowest probability items from the list, which are the worst performing on the unigram LM.
II-B Attention
Attention, particularly selective attention, has been widely studied under perception, psychophysics, and psychology. Selective attention can be conceived as “the programming by the O of which stimuli will be processed or encoded and in what order this will occur” [60]. While this definition has its roots in visual perception, it has uncanny similarities with the recently formulated attention [61, 62] (which stimuli will be processed) and positional encoding (in what order this will occur) [62] in LLMs. We discuss both in sections II-C and II-D, respectively.
@@ II-C Attention in LLMs
The attention mechanism computes a representation of the input sequences by relating different positions (tokens) of these sequences. There are various approaches to calculating and implementing attention, out of which some famous types are given below.
II-C1 Self-Attention [62]
The self-attention is also known as intra-attention since all the queries, keys, and values come from the same block (encoder or decoder). The self-attention layer connects all the sequence positions with $O(1)$ space complexity which is highly desirable for learning long-range dependencies in the input.
II-C2 Cross Attention
In encoder-decoder architectures, the outputs of the encoder blocks act as the queries to the intermediate representation of the decoder, which provides the keys and values to calculate a representation of the decoder conditioned on the encoder. This attention is called cross-attention.
II-C3 Full Attention
The naive implementation of calculating self-attention is known as full attention.
II-C4 Sparse Attention [63]
The self-attention has a time complexity of $O(n^{2})$, which becomes prohibitive when scaling the LLMs to large context windows. An approximation to the self-attention was proposed in [63], which greatly enhanced the capacity of GPT series LLMs to process a greater number of input tokens in a reasonable time.
II-C5 Flash Attention [64]
The bottleneck for calculating the attention using GPUs lies in the memory access rather than the computational speed. Flash Attention uses the classical input tiling approach to process the blocks of the input in GPU on-chip SRAM rather than doing IO for every token from the High Bandwith Memory (HBM). An extension of this approach to sparse attention follows the speed gains of the full attention implementation. This trick allows even greater context-length windows in the LLMs as compared to those LLMs with sparse attention.
II-D Encoding Positions
The attention modules do not consider the order of processing by design. Transformer [62] introduced “positional encodings” to feed information about the position of the tokens in input sequences. Several variants of positional encoding have been proposed [65, 66]. Interestingly, a recent study [67] suggests that adding this information may not matter for the state-of-the-art decoder-only Transformers.
II-D1 Absolute
This is the most straightforward approach to adding the sequence order information by assigning a unique identifier to each position of the sequence before passing it to the attention module.
II-D2 Relative
To pass the information on the relative dependencies of different tokens appearing at different locations in the sequence, a relative positional encoding is calculated by some kind of learning. Two famous types of relative encodings are:
Alibi: [65] In this approach, a scalar bias is subtracted from the attention score calculated using two tokens which increases with the distance between the positions of the tokens. This learned approach effectively favors using recent tokens for attention.
RoPE: Keys, queries, and values are all vectors in the LLMs. RoPE [66] involves the rotation of the query and key representations at an angle proportional to their absolute positions of the tokens in the input sequence. This step results in a relative positional encoding scheme which decays with the distance between the tokens.
II-E Activation Functions
The activation functions serve a crucial role in the curve-fitting abilities of the neural networks, as proved in [68]. The modern activation functions used in LLMs are different from the earlier squashing functions but are critical to the success of LLMs. We discuss these activation functions in this section.
II-E1 ReLU [69]
Rectified linear unit (ReLU) is defined as
@@ II-E1 ReLU [69]
II-E2 GeLU [70]
Gaussian Error Linear Unit (GeLU) is the combination of ReLU, dropout [71] and zoneout [72]. It is the most widely used activation function in contemporary LLM literature.
II-E3 GLU variants [73]
Gated Linear Unit [74] is a neural network layer that is an element-wise product ($\otimes$) of a linear transformation and a sigmoid transformed ($\sigma$) linear projection of the input given as
@@ II-E3 GLU variants [73]
GLU was modified in [73] to evaluate the effect of different variations in the training and testing of transformers, resulting in better empirical results. Here are the different GLU variations introduced in [73] and used in LLMs.
@@
II-F Layer Normalization
Layer normalization leads to faster convergence and is a widely used component in transformers. In this section, we provide different normalization techniques widely used in LLM literature.
II-F1 LayerNorm
Layer norm computes statistics over all the hidden units in a layer $(l)$ as follows:
$u^{l}=\frac{1}{n}\sum_{i}^{n}a_{i}^{l}\hskip 20.00003pt\sigma^{l}=\sqrt{\frac{1}{n}\sum_{i}^{n}(a_{i}^{l}-u^{l})^{2}},$ (3)
where $n$ is the number of neurons in the layer $l$ and $a_{i}^{l}$ is the summed input of the $i$ neuron in layer $l$. LayerNorm provides invariance to rescaling of the weights and re-centering of the distribution.
II-F2 RMSNorm
[75] proposed that the invariance properties of LayerNorm are spurious, and we can achieve the same performance benefits as we get from LayerNorm by using a computationally efficient normalization technique that trades off re-centering invariance with speed. LayerNorm gives the normalized summed input to layer $l$ as follows
$\overline{a_{i}^{l}}=\frac{a_{i}^{l}-u^{l}}{\sigma}g_{i}^{l}$ (4)
where $g_{i}^{l}$ is the gain parameter. RMSNorm [75] modifies $\overline{a_{i}^{l}}$ as
$\overline{a_{i}^{l}}=\frac{a_{i}^{l}}{\textnormal{RMS}(\mathbf{a}^{l})}g_{i}^{l},\hskip 3.00003pt\textnormal{where}\hskip 3.00003pt\textnormal{RMS}(\mathbf{a}^{l})=\sqrt{\frac{1}{n}\sum_{i}^{n}(a_{i}^{l})^{2}}.$ (5)
II-F3 Pre-Norm and Post-Norm
LLMs use transformer [62] architecture with some variations. The original implementation [62] used layer normalization after the residual connection, commonly called post-LN, concerning the order of Multihead attention – Residual – LN. There is another order of the normalization, referred to as pre-LN [76] due to the position of the normalization step before the self-attention layer as in LN – Multihead attention – Residual. Pre-LN is known to provide more stability in the training [77].
II-F4 DeepNorm
While pre-LN has certain benefits over post-LN training, pre-LN training has an unwanted effect on the gradients [77]. The earlier layers have larger gradients than those at the bottom. DeepNorm [78] mitigates these adverse effects on the gradients. It is given as
$\mathbf{x}^{l_{f}}=LN(\alpha\mathbf{x}^{l_{p}}+G^{l_{p}}(\mathbf{x}^{l_{p}},{\boldmath\theta}{{}^{l_{p}}}),$ (6)
where $\alpha$ is a constant and $\theta^{l_{p}}$ represents the parameters of layer $l_{p}$. These parameters are scaled by another constant $\beta$. Both of these constants depend only on the architecture.
II-G Distributed LLM Training
@@
II-G1 Data Parallelism
Data parallelism replicates the model on multiple devices where data in a batch gets divided across devices. At the end of each training iteration weights are synchronized across all devices.
II-G2 Tensor Parallelism
Tensor parallelism shards a tensor computation across devices. It is also known as horizontal parallelism or intra-layer model parallelism.
II-G3 Pipeline Parallelism
Pipeline parallelism shards model layers across different devices. This is also known as vertical parallelism.
II-G4 Model Parallelism
A combination of tensor and pipeline parallelism is known as model parallelism.
II-G5 3D Parallelism
A combination of data, tensor, and model parallelism is known as 3D parallelism.
II-G6 Optimizer Parallelism
Optimizer parallelism also known as zero redundancy optimizer [37] implements optimizer state partitioning, gradient partitioning, and parameter partitioning across devices to reduce memory consumption while keeping the communication costs as low as possible.
II-H Libraries
@@ II-I Data PreProcessing
II-I1 Quality Filtering
For better results, training data quality is essential. Some approaches to filtering data are: 1) classifier-based and 2) heuristics-based. Classifier-based approaches train a classifier on high-quality data and predict the quality of text for filtering, whereas heuristics-based employ some rules for filtering like language, metrics, statistics, and keywords.
II-I2 Data Deduplication
Duplicated data can affect model performance and increase data memorization; therefore, to train LLMs, data deduplication is one of the preprocessing steps. This can be performed at multiple levels, like sentences, documents, and datasets.
II-I3 Privacy Reduction
Most of the training data for LLMs is collected through web sources. This data contains private information; therefore, many LLMs employ heuristics-based methods to filter information such as names, addresses, and phone numbers to avoid learning personal information.
II-J Architectures
Here we discuss the variants of the transformer architectures at a higher level which arise due to the difference in the application of the attention and the connection of transformer blocks. An illustration of attention patterns of these architectures is shown in Figure 4.
II-J1 Encoder Decoder
Transformers were originally designed as sequence transduction models and followed other prevalent model architectures for machine translation systems. They selected encoder-decoder architecture to train human language translation tasks. This architecture is adopted by [10, 89]. In this architectural scheme, an encoder encodes the input sequences to variable length context vectors, which are then passed to the decoder to maximize a joint objective of minimizing the gap between predicted token labels and the actual target token labels.
II-J2 Causal Decoder
The underlying objective of an LLM is to predict the next token based on the input sequence. While additional information from the encoder binds the prediction strongly to the context, it is found in practice that the LLMs can perform well in the absence of encoder [90], relying only on the decoder. Similar to the original encoder-decoder architecture’s decoder block, this decoder restricts the flow of information backward, i.e., the predicted token $t_{k}$ only depends on the tokens preceded by and up to $t_{k-1}$. This is the most widely used variant in the state-of-the-art LLMs.
II-J3 Prefix Decoder
The causal masked attention is reasonable in the encoder-decoder architectures where the encoder can attend to all the tokens in the sentence from every position using self-attention. This means that the encoder can also attend to tokens $t_{k+1}$ to $t_{n}$ in addition to the tokens from $t_{1}$ to $t_{k-1}$ while calculating the representation for $t_{k}$. But when we drop the encoder and only keep the decoder, we also lose this flexibility in attention. A variation in the decoder-only architectures is by changing the mask from strictly causal to fully visible on a portion of the input sequence, as shown in Figure 4. The Prefix decoder is also known as non-causal decoder architecture.
@@ II-K Pre-Training Objectives
II-K1 Full Language Modeling
An autoregressive language modeling objective where the model is asked to predict future tokens given the previous tokens, an example is shown in Figure 5.
II-K2 Prefix Language Modeling
A non-causal training objective, where a prefix is chosen randomly and only remaining target tokens are used to calculate the loss. An example is shown in Figure 5.
II-K3 Masked Language Modeling
In this training objective, tokens or spans (a sequence of tokens) are masked randomly and the model is asked to predict masked tokens given the past and future context. An example is shown in Figure 5.
II-K4 Unified Language Modeling
Unified language modeling [92] is a combination of causal, non-causal, and masked language training objectives. Here in masked language modeling, the attention is not bidirectional but unidirectional, attending either left-to-right or right-to-left context.
II-L Model Adaptation
@@
II-L1 Pre-Training
In the very first stage, the model is trained in a self-supervised manner on a large corpus to predict the next tokens given the input. The design choices of LLMs vary from encoder-decoder to decoder-only architectures with different building blocks and loss functions in sections II-FII-E,  II-K.
II-L2 Fine-Tuning
@@ III-A Pre-Trained LLMs
Here, we provide summaries of various well-known pre-trained LLMs with significant discoveries, changing the course of research and development in NLP. These LLMs have considerably improved the performance in NLU and NLG domains, and are widely fine-tuned for downstream tasks.
@@ III-C Increasing Context Window
III-D Robotics
LLMs have been rapidly adopted across various domains in the scientific community due to their multipurpose capabilities [46]. In robotics research, the LLMs have very promising applications as well, such as enhancing human-robot interaction [28, 182, 183, 184], task planning [185, 186, 187], navigation [188, 189], and learning [190, 191]. They can enable robots to understand and generate natural language, aiding in instruction following, data annotation, and collaborative problem-solving. They can facilitate continuous learning by allowing robots to access and integrate information from a wide range of sources. This can help robots acquire new skills, adapt to changes, and refine their performance based on real-time data.
LLMs have also started assisting in simulating environments for testing and offer potential for innovative research in robotics, despite challenges like bias mitigation and integration complexity. The work in [192] focuses on personalizing robot household cleanup tasks. By combining language-based planning and perception with LLMs, such that having users provide object placement examples, which the LLM summarizes to generate generalized preferences, they show that robots can generalize user preferences from a few examples. An embodied LLM is introduced in [26], which employs a Transformer-based language model where sensor inputs are embedded alongside language tokens, enabling joint processing to enhance decision-making in real-world scenarios. The model is trained end-to-end for various embodied tasks, achieving positive transfer from diverse training across language and vision domains. LLMs have also been explored as zero-shot human models for enhancing human-robot interaction.
The study in [28] demonstrates that LLMs, trained on vast text data, can serve as effective human models for certain HRI tasks, achieving predictive performance comparable to specialized machine-learning models. However, limitations were identified, such as sensitivity to prompts and difficulties with spatial/numerical reasoning. In another study [193], the authors enable LLMs to reason over sources of natural language feedback, forming an “inner monologue” that enhances their ability to process and plan actions in robotic control scenarios. They combine LLMs with various forms of textual feedback, allowing the LLMs to incorporate conclusions into their decision-making process for improving the execution of user instructions in different domains, including simulated and real-world robotic tasks involving tabletop rearrangement and mobile manipulation. All of these studies employ LLMs as the core mechanism for assimilating everyday intuitive knowledge into the functionality of robotic systems.
Planning: LLMs are increasingly integral in robotics, particularly for strategic planning [185, 194, 195]. Their proficiency in processing and generating natural language is crucial for enhancing human-robot interaction and enabling robots to understand and execute complex tasks based on verbal instructions. LLMs also play a key role in task planning, a higher-level cognitive process involving the determination of sequential actions needed to achieve specific goals. This proficiency is crucial across a spectrum of applications, from autonomous manufacturing processes to household chores, where the ability to understand and execute multi-step instructions is of paramount significance.
Manipulation: In the area of manipulation [196, 197, 198, 199], LLMs enhance a robot’s dexterity and adaptability, excelling in tasks like object recognition, grasping, and collaboration. They analyze visual and spatial information to determine the most effective approach to interact with objects, proving invaluable in operations requiring precision and flexibility, such as surgical procedures or assembly line tasks. They also enable the integration of sensor inputs and linguistic cues in an embodied framework, enhancing decision-making in real-world scenarios. It enhances the model’s performance across various embodied tasks by allowing it to gather insights and generalize from diverse training data spanning language and vision domains.
Navigation: LLMs have revolutionized the navigation in robotics [200, 201, 202, 203], offering significant potential to enhance a robot’s ability to navigate complex environments with precision and adaptability. Motion planning [188], in particular, stands out as a critical domain where LLMs have shown remarkable promise, excelling in generating feasible paths and trajectories for robots, accounting for intricate environmental details. This ability proves particularly valuable in scenarios requiring precise and dynamically adaptable navigation, as observed in environments like warehouses, transport and healthcare facilities, and smart residences. LLMs have also played a key role in localization and mapping, which are foundational components for successful robot navigation. They empower robots to determine their precise position within an environment while concurrently constructing or updating a spatial representation of their surroundings. This capability is crucial for tasks demanding spatial awareness, including autonomous exploration, search and rescue missions, and the operations of mobile robots. They have also contributed significantly to the proficiency of collision-free navigation within the environment while accounting for obstacles and dynamic alterations, playing an important role in scenarios where robots are tasked with traversing predefined paths with accuracy and reliability, as seen in the operations of automated guided vehicles (AGVs) and delivery robots (e.g., SADRs – pedestrian sized robots that deliver items to customers without the involvement of a delivery person).
III-E Multimodal LLMs
III-F Augmented LLMs

### 20 Feb 2024 v8

@@ Abstract
Figure 1: The trend of papers released over years containing keywords “Large Language Model”, “Large Language Model + Fine-Tuning”, and “Large Language Model + Alignment”.
@@ I Introduction
In addition to better generalization and domain adaptation, LLMs appear to have emergent abilities, such as reasoning, planning, decision-making, in-context learning, answering in zero-shot settings, etc. These abilities are known to be acquired by them due to their gigantic scale even when the pre-trained LLMs are not trained specifically to possess these attributes [22, 23, 24]. Such abilities have led LLMs widely adopted in diverse settings including, multi-modal, robotics, tool manipulation, question answering, autonomous agents, etc. Various improvements have also been suggested in these areas either by task-specific training [25, 26, 27, 28, 29, 30, 31] or better prompting [32].
Figure 3: A broader overview of LLMs, dividing LLMs into seven branches: 1. Pre-Training 2. Fine-Tuning 3. Efficient 4. Inference 5. Evaluation 6. Applications 7. Challenges
The LLMs abilities to solve diverse tasks with human-level performance come at a cost of slow training and inference, extensive hardware requirements, and higher running costs. Such requirements have limited their adoption and opened up opportunities to devise better architectures [15, 33, 34, 35] and training strategies [36, 37, 21, 38, 39, 40, 41]. Parameter efficient tuning [38, 41, 40], pruning [42, 43], quantization [44, 45], knowledge distillation, and context length interpolation [46, 47, 48, 49] among others are some of the methods widely studied for efficient LLM utilization.
@@
  • We present a survey on the developments in LLM research providing a concise comprehensive overview of the direction.
@@
  • We summarize major findings of the popular contributions and provide a detailed discussion on the key design and development aspects of LLMs to help practitioners effectively leverage this technology.
  • In this self-contained article, we cover a range of concepts to present the general direction of LLMs comprehensively, including background, pre-training, fine-tuning, multi-modal LLMs, augmented LLMs, LLMs-powered agents, datasets, evaluation, etc.
We loosely follow the existing terminologies to ensure a standardized outlook of this research direction. For instance, following [50], our survey discusses pre-trained LLMs with 10B parameters or more. We refer the readers interested in smaller pre-trained models to [51, 52, 53].
The organization of this paper is as follows. Section II discusses the background of LLMs. Section III focuses on LLMs overview, architectures, training pipelines and strategies, fine-tuning, and utilization in different aspects. Section IV highlights the configuration and parameters that play a crucial role in the functioning of these models. Summary and discussions are presented in section III-H. The LLM training and evaluation, datasets, and benchmarks are discussed in section V, followed by challenges and future directions and conclusion in sections VII and VIII, respectively.
@@ II Background
We provide the relevant background to understand the fundamentals related to LLMs in this section. We briefly discuss necessary components in LLMs and refer the readers interested in details to the original works.
@@ II-A Tokenization
Tokenization [59] is an essential pre-processing step in LLM training that parses the text into non-decomposing units called tokens. Tokens can be characters, subwords [60], symbols [61], or words, depending on the tokenization process. Some of the commonly used tokenization schemes in LLMs include wordpiece [62], byte pair encoding (BPE) [61], and unigramLM [60]. Readers are encouraged to refer to [63] for a detailed survey.
II-B Encoding Positions
The transformer processes input sequences in parallel and independently of each other. Moreover, the attention module in the transformer does not capture positional information. As a result, positional encodings were introduced in transformer [64], where a positional embedding vector is added with the token embedding. Variants of positional embedding include absolute, relative, or learned positional encodings. Within relative encoding, Alibi and RoPE are two widely used positional embeddings in LLMs.
Alibi [65]: It subtracts a scalar bias from the attention score that increases with the distance between token positions. This favors using recent tokens for attention.
RoPE [66]: It rotates query and key representations at an angle proportional to the token absolute position in the input sequence, resulting in a relative positional encoding scheme which decays with the distance between the tokens.
@@ II-C Attention in LLMs
The attention assigns weights to input tokens based on importance so that the model emphasizes more on relevant over irrelevant tokens. Attention in transformers [64] calculates query, key, and value mappings for input sequences, where the attention score is obtained by multiplying the query and key, later used to weight values. We discuss different attention strategies used in LLMs below.
Self-Attention [64]: Calculates attention using queries, keys, and values from the same block (encoder or decoder).
Cross Attention: It is used in encoder-decoder architecture, where encoder outputs are the queries, and key-value pairs come from the decoder.
Sparse Attention [67]: Self-attention has $O(n^{2})$ time complexity which becomes infeasible for large sequences. To speed up the computation, sparse attention [67] iteratively calculates the attention in sliding windows for speed gains.
Flash Attention [68]: Memory access is the major bottleneck in calculating attention using GPUs. To speed up, flash attention employs input tiling to minimize the memory reads and writes between the GPU high bandwidth memory (HBM) and the on-chip SRAM.
II-D Activation Functions
The activation functions serve a crucial role in the curve-fitting abilities of the neural networks [69]. We discuss activation functions used in LLMs in this section.
ReLU [70]: Rectified linear unit (ReLU) is defined as
@@ II-D Activation Functions
GeLU [71]: Gaussian Error Linear Unit (GeLU) is the combination of ReLU, dropout [72] and zoneout [73].
GLU variants [74]: Gated Linear Unit [75] is a neural network layer that is an element-wise product ($\otimes$) of a linear transformation and a sigmoid transformed ($\sigma$) linear projection of the input given as
@@
Other GLU variants [74] used in LLMs are:
@@
II-E Layer Normalization
Layer normalization leads to faster convergence and is an integrated component of transformers [64]. Besides LayerNorm [76] and RMSNorm [77], LLMs use pre-layer normalization [78], applying it before multi-head attention (MHA). Pre-norm is shown to provide training stability in LLMs. Another normalization variant, DeepNorm [79] fixes larger gradient issue in pre-norm.
II-F Distributed LLM Training
@@
Data Parallelism: Data parallelism replicates the model on multiple devices where data in a batch gets divided across devices. At the end of each training iteration weights are synchronized across all devices.
Tensor Parallelism: Tensor parallelism shards a tensor computation across devices. It is also known as horizontal parallelism or intra-layer model parallelism.
Pipeline Parallelism: Pipeline parallelism shards model layers across different devices. This is also known as vertical parallelism.
Model Parallelism: A combination of tensor and pipeline parallelism is known as model parallelism.
3D Parallelism: A combination of data, tensor, and model parallelism is known as 3D parallelism.
Optimizer Parallelism: Optimizer parallelism also known as zero redundancy optimizer [37] implements optimizer state partitioning, gradient partitioning, and parameter partitioning across devices to reduce memory consumption while keeping the communication costs as low as possible.
II-G Libraries
@@ II-H Data PreProcessing
Quality Filtering: For better results, training data quality is essential. Some approaches to filtering data are: 1) classifier-based and 2) heuristics-based. Classifier-based approaches train a classifier on high-quality data and predict the quality of text for filtering, whereas heuristics-based employ some rules for filtering like language, metrics, statistics, and keywords.
Data Deduplication: Duplicated data can affect model performance and increase data memorization; therefore, to train LLMs, data deduplication is one of the preprocessing steps. This can be performed at multiple levels, like sentences, documents, and datasets.
Privacy Reduction: Most of the training data for LLMs is collected through web sources. This data contains private information; therefore, many LLMs employ heuristics-based methods to filter information such as names, addresses, and phone numbers to avoid learning personal information.
II-I Architectures
Here we discuss the variants of the transformer architectures used in LLMs. The difference arises due to the application of the attention and the connection of transformer blocks. An illustration of attention patterns of these architectures is shown in Figure 4.
Encoder Decoder: This architecture processes inputs through the encoder and passes the intermediate representation to the decoder to generate the output. Here, the encoder sees the complete sequence utilizing self-attention whereas the decoder processes the sequence one after other with implementing cross-attention.
Causal Decoder: A type of architecture that does not have encoder and processes and generates output using decoder, where predicted token depends only on the previous time steps.
Prefix Decoder: Also known as non-causal decoder, where the attention calculation is not strictly dependant on the past information and the attention is bidirectional. An example of non-causal attention mask is shown in Figure 4.
@@ II-J Pre-Training Objectives
Full Language Modeling: An autoregressive language modeling objective where the model is asked to predict future tokens given the previous tokens, an example is shown in Figure 5.
Prefix Language Modeling: A non-causal training objective, where a prefix is chosen randomly and only remaining target tokens are used to calculate the loss. An example is shown in Figure 5.
Masked Language Modeling: In this training objective, tokens or spans (a sequence of tokens) are masked randomly and the model is asked to predict masked tokens given the past and future context. An example is shown in Figure 5.
Unified Language Modeling: Unified language modeling [91] is a combination of causal, non-causal, and masked language training objectives. Here in masked language modeling, the attention is not bidirectional but unidirectional, attending either left-to-right or right-to-left context.
II-K LLMs Adaptation Stages
@@
II-K1 Pre-Training
In the very first stage, the model is trained in a self-supervised manner on a large corpus to predict the next tokens given the input. The design choices of LLMs vary from encoder-decoder to decoder-only architectures with different building blocks and loss functions in sections II-EII-D,  II-J.
II-K2 Fine-Tuning
@@ III-A Pre-Trained LLMs
Here, we provide summaries of various well-known pre-trained LLMs with significant discoveries, changing the course of research and development in NLP. These LLMs have considerably improved the performance in NLU and NLG domains, and are widely fine-tuned for downstream tasks. We identify critical information of pre-trained LLMs in Table I and II that improve their performance.
@@ III-C Increasing Context Window
III-D Augmented LLMs